OpenStudy (anonymous):

ideal gas law PV=nRT what physical quantity is represented on each side of the equation?

OpenStudy (rushwr):

P = Pressure in pascal V= Volume in cubic meters n= no. of moles R= Universal gas constant T= temperature in Kelvin

OpenStudy (anonymous):

This is a perfect example of the advantage of dimensional analysis... its a statement about energy: \[P = \frac{Force}{Area} = \frac{Force}{Length^{2}} \ \ \ \& \ \ \ V=Length^{3}\] So: \[PV = \frac{Force}{Length^{2}}*Length^{3}=Force*Length=Energy\] The same goes for the other side of the equation: R is the universal gas constant that has units of Energy per (Temperature*mole)... so multiplied by the number of moles and the Temperature gives you energy